# DyGMamba: Efficiently Modeling Long-Term Temporal Dependency on Continuous-Time Dynamic Graphs with State Space Models

### v1

@@ 3.2 Learning from Edge-Specific Temporal Patterns
To capture edge-specific temporal patterns, we use another time-level SSM block consisting of $l_{T}$ layers. We first find out $k$ temporally nearest historical interactions between $u$ and $v$ before $t$ and sort them in the chronological order, i.e., $\{(u,v,t_{0}),...,(u,v,t_{k-1})|t_{0}<...<t_{k-1},t_{k-1}<t\}$. Then we construct a timestamp sequence $\{t_{0},t_{1},...,t_{k-1},t\}$ based on these interactions and the prediction timestamp $t$. We compute the time difference between each neighboring pair of them and further get a time difference sequence $\{t_{1}-t_{0},t_{2}-t_{1},...,t-t_{k-1}\}$, representing the change of time interval between two identical interactions. Each element in this sequence is input into the time encoding function stated above to get a edge-specific (specific to the edge $(u,v,t)$) time feature. The features are stacked into a feature matrix $\mathbf{H}_{u,v}^{t}\in\mathbb{R}^{k\times d_{T}}$ and mapped by an MLP $f_{\text{map}1}(\cdot):\mathbb{R}^{d_{T}}\to\mathbb{R}^{\gamma d}$ ($\gamma\in[0,1]$ is a hyperparameter), i.e., $\mathbf{H}_{u,v}^{t}:=f_{\text{map}1}(\mathbf{H}_{u,v}^{t})$. A time-level SSM layer takes $\mathbf{H}_{u,v}^{t}$ as input and computes
@@
After the time-level SSM block, we compute a compressed representation to represent the edge-specific temporal pattern
$\mathbf{h}_{u,v}^{t}=\text{MeanPooling}(\mathbf{H}_{u,v}^{t}),$ (7)
where $\text{MeanPooling}(\cdot)$ averages over $k$ interactions. As a result, we have $\mathbf{h}_{u,v}^{t}\in\mathbb{R}^{\gamma d}$ to represent the temporal pattern specific to the edge $(u,v,t)$. To leverage learned temporal pattern, we use it to dynamically select the information from node representations $\mathbf{h}_{u}^{t}$ and $\mathbf{h}_{v}^{t}$
$\tilde{\mathbf{h}}_{u,v}^{t}=f_{\text{map}2}(\mathbf{h}_{u,v}^{t})\in\mathbb{R}^{4d};$ (8a)
$\tilde{\mathbf{h}}_{\theta}^{t}={\mathbf{w}_{\text{agg}}}^{\top}\mathbf{H}_{\theta}^{t}\in\mathbb{R}^{4d},\ \text{where}\ \mathbf{w}_{\text{agg}}=f_{\text{map}3}(\mathbf{H}_{\theta}^{t})\in\mathbb{R}^{|\textit{Nei}_{\theta}^{t}|+1};$ (8b)
$\alpha_{u}=f^{\prime}(\tilde{\mathbf{h}}_{v}^{t})*\tilde{\mathbf{h}}_{u,v}^{t}\in\mathbb{R}^{4d},\ \alpha_{v}=f^{\prime}(\tilde{\mathbf{h}}_{u}^{t})*\tilde{\mathbf{h}}_{u,v}^{t}\in\mathbb{R}^{4d};$ (8c)
$\mathbf{h}_{\theta}^{t}={\beta_{\theta}}^{\top}\mathbf{H}_{\theta}^{t},\ \text{where}\ \beta_{\theta}=\text{Softmax}(\mathbf{H}_{\theta}^{t}\alpha_{\theta})\in\mathbb{R}^{|\textit{Nei}_{\theta}^{t}|+1}.$ (8d)
$f_{\text{map}2}(\cdot):\mathbb{R}^{\gamma d}\to\mathbb{R}^{4d}$ and $f_{\text{map}3}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{1}$ are two mapping MLPs. $f_{\text{agg}}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$ is another MLP introducing training parameters. $\alpha_{\theta}$ can be viewed as a key used to select important historical information from the node $\theta$ (either $u$ or $v$) at $t$. Note that $\alpha_{u}$/$\alpha_{v}$ is computed by considering both the edge-specific temporal pattern and the opposite node $v$/$u$. In the node-level SSM block, we separately model the one-hop temporal neighbors of each node $\theta$, making it hard to connect $u$ and $v$. Computing $\alpha_{\theta}$ as Eq. 8c helps to strengthen the connection between both nodes and meanwhile incorporates the learned temporal pattern. $\beta_{\theta}$ is derived by transforming the queried results based on $\alpha_{\theta}$ into weights. It is then used to compute a weighted-sum of all temporal neighbors for representing $\theta$ at $t$, i.e., $\mathbf{h}_{\theta}^{t}$. Finally, we output the representations of $u$, $v$ and the edge-specific temporal pattern by employing two output MLPs $f_{\text{out}1}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{d_{N}}$ and $f_{\text{out}2}(\cdot):\mathbb{R}^{\gamma d}\to\mathbb{R}^{d_{N}}$
@@ 3.3 Leveraging Learned Representations for Link Prediction
We leverage $\mathbf{h}_{\theta}^{t}$ and $\mathbf{h}_{u,v}^{t}$ for dynamic link prediction. We employ a prediction MLP, i.e., $f_{\text{LP}}(\cdot):\mathbb{R}^{3d_{N}}\to\mathbb{R}$, as the predictor. The probability of existing a link $(u,v,t)$ is computed as $\hat{y}(u,v,t)=\text{Sigmoid}(f_{\text{LP}}(\mathbf{h}_{u}^{t}\|\mathbf{h}_{v}^{t}\|\mathbf{h}_{u,v}^{t}))$. For model parameter learning, we use the following loss function
$\mathcal{L}=-\frac{1}{2M}\sum_{2M}\left(y(u,v,t)\log(\hat{y}(u,v,t))+(1-y(u,v,t))\log(1-\hat{y}(u,v,t))\right).$ (10)
@@ 4 Experiments
In Sec. 4.2, we validate DyGMamba’s ability in CTDG representation learning by comparing it with baseline methods on dynamic link prediction [^2] To supplement, we also validate on the dynamic node classification task. Since current mainstream datasets of this task requires no long-term temporal reasoning, we put the discussion in App. G, and present ablation studies to demonstrate the effectiveness of model components. In Sec. 4.3, we further show DyGMamba’s efficiency in terms of model size, training time and GPU memory consumption. We also show that DyGMamba achieves much stronger scalability in modeling long-term temporal information compared with DyGFormer.
@@ 4.1 Experimental Setting
We consider seven real-world CTDG datasets collected by (Poursafaei et al., 2022), i.e., LastFM, Enron, MOOC, Reddit, Wikipedia, UCI and Social Evo.. Dataset statistics are presented in App. B. Among them, we take LastFM, Enron and MOOC as long-range temporal dependent datasets because according to (Yu et al., 2023), much longer histories are needed for optimal representation learning on them. We compare DyGMamba with ten recent CTDG baseline models, i.e., JODIE (Kumar et al., 2019), DyRep (Trivedi et al., 2019), TGAT (Xu et al., 2020), TGN (Rossi et al., 2020), CAWN (Wang et al., 2021c), EdgeBank (Poursafaei et al., 2022), TCL (Wang et al., 2021a), GraphMixer (Cong et al., 2023), DyGFormer (Yu et al., 2023) and CTAN (Gravina et al., 2024). Among them, only DyGFormer and CTAN are designed for long-range temporal information propagation. Detailed descriptions of baseline methods are presented in App. C. We also implemented FreeDyG (Tian et al., 2024) by using its official code repository, however, on LastFM, we find that FreeDyG’s loss cannot converge and the reported results are not reproducible. So we do not report its performance in our paper.
Evaluation Settings and Metrics.
We employ two evaluation settings following previous works: the transductive and inductive settings. In the transductive setting, models are asked to predict the future links between the nodes that have been seen during training, while in the inductive setting, models should predict the links between previously unseen nodes. As suggested in (Poursafaei et al., 2022), we do link prediction evaluation using three negative sampling strategies (NSSs): random, historical and inductive. Historical NSS is only considered under the transductive setting. See App. E for detailed explanations. We employ two metrics, i.e., average precision (AP) and area under the receiver operating characteristic curve (AUC-ROC) We chronologically split each dataset with the ratio of 70%/15%/15% for training/validation/testing.
Implementation Details.
[…]
@@
4.2 Performance Analysis
Comparative Study on Benchmark Datasets.
We report the AP of transductive and inductive link prediction in Table 1 and 2 (AUC-ROC reported in Table 8 and 9 in App. F). We find that: (1) DyGMamba constantly ranks top $1$ under the random NSS, showing a superior performance; (2) Under the historical and inductive NSS, DyGMamba can achieve the best average rank compared with all baselines. More importantly, it performs much better on the datasets where encoding longer-term temporal dependencies is necessary, e.g., on LastFM, Enron and MOOC which requires larger $\rho$. This shows DyGMamba’s ability in modeling long-term temporal dependent CTDGs; (3) Among the models that can do long range propagation of information over time (i.e., DyGFormer, CTAN and DyGMamba), DyGMamba achieves the best average rank under any NSS setting in both transductive and inductive link prediction. On the long-range temporal datasets LastFM, Enron and MOOC, DyGMamba outperforms DyGFormer and CTAN in most cases; (4) CTAN achieves much better results in transductive than in inductive link prediction. This is because CTAN requires multi-hop temporal neighbors to learn node representations, which is difficult for unseen nodes. By contrast, DyGMamba and DyGFormer requires only one-hop temporal neighbors, thus performing much better in inductive link prediction.
@@ 4.2 Performance Analysis
Ablation Study.
We conduct two ablation studies to study the effectiveness of model components. In study A, we make a model variant (Varaint A) by removing the time-level SSM block and restrain our model from learning temporal patterns (information selection is substituted by mean pooling over the output of Eq. 5). In study B, we make a model variant (Variant B) by removing the Mamba SSM layers (Eq. 4) in the node-level SSM block. We have three findings: (1) From Table 3 and 4, we observe that under any NSS setting, Variant A is constantly beaten by the complete model in both transductive and inductive link prediction. This proves the effectiveness of learning edge-specific temporal patterns; (2) The performance gap between Variant A and DyGMamba enlarges as model considers more temporal neighbors, especially under the historical and inductive NSS (e.g., Enron vs. Social Evo.). This implies that dynamic information selection with temporal patterns is more contributive when more temporal information is introduced; (3) We also witness a performance drop from Table 3 and 4 if we switch DyGMamba to Variant B. This indicates the importance of encoding the one-hop temporal neighbors with SSM layers for capturing graph dynamics.
Table 3: Ablation studies under transductive setting. R/I means random/inductive NSS.
@@
Table 4: Ablation studies under inductive setting. R/I means random/inductive NSS.
@@ 4.3 Efficiency Analysis
Figure 2: Efficiency comparison on four datasets among DyGMamba and five baselines in terms of number (#) of parameters, training time per epoch and GPU memory. The performance metric here is the AP of transductive link prediction under random NSS.
@@
Fig. 2 compares DyGMamba with five baselines in terms of number of parameters (model size), per epoch training time and GPU memory consumption during training [^3] The baselines not included here are either extremely inefficient (e.g., CAWN) or inferior in performance (e.g., DyRep). Complete statistics of all baseline models presented in App. H. We find that: (1) DyGMamba uses very few parameters while maintaining the best performance, showing a strong parameter efficiency. Only CTAN constantly uses fewer parameters than DyGMamba, however, its performance is much worse; (2) DyGMamba is much more efficient than DyGFormer. Note that the lengths of the sequences input into the SSM and the Transformer are the same ($\rho/p=32$), which ensures a fair comparison; (3) Although DyGMamba generally consumes more GPU memory and per epoch training time compared with most baselines, the gap of consumption is not very large in most cases. To model more temporal neighbors for long-range temporal dependent datasets, DyGMamba naturally requires more computational resources, thus enlarging the consumption gap. DyGFormer shows the same trend as DyGMamba since it also captures long-term temporal dependencies by considering more temporal neighbors; (4) CTAN requires very few computational resources. However, on long-range temporal dependent datasets, it is beaten by DyGFormer and DyGMamba with a great margin. Besides, CTAN is also hard to converge. Although it requires very little per epoch training time, it requires much more epochs to reach the best performance, leading to a long total training time. See App. H.2 and H.3 for details.
Impact of Patch Size on Scalability.
@@
Patching treats $p$ temporal neighbors as one patch and thus decreases the sequence length by $p$ times. This is very helpful in cutting the consumption of GPU memory and training/evaluation time. However, patching introduces excessive parameters because it is done through $f_{N}$, $f_{E}$, $f_{T}$ and $f_{F}$ (Eq. 3) whose size increases as the patch size grows. Fig. 3b shows the numbers of parameters of DyGFormer and DyGmamba with different patch sizes on a long-term temporal dependent dataset Enron. We find that patching greatly affects model sizes. To further study how patching affects DyGMamba, we decrease the patch size gradually from $8$ to $1$ and track DyGMamba’s performance (Fig. 3a) as well as efficiency (Fig. 3b to 3d) on Enron. Meanwhile, we also keep track on DyGFormer under the same patch size for comparison. We have several findings: (1) Whatever the patch size is, DyGMamba always consumes fewer parameters, less GPU memory and per epoch training time, showing its high efficiency; (2) While both models require increasing computational budgets as the patch size decreases, the speed of increase is much lower for DyGMamba, demonstrating its strong scalability in modeling longer sequences; (3) Different trends in performance change are observed between two models. While DyGFormer performs worse, DyGMamba can benefit from a smaller patch size, indicating its strong ability to capture nuanced temporal details even if the sequence becomes much longer. Note that the models use fewer parameters under smaller patch size. This also shows that DyGMamba can achieve much stronger parameter efficiency by reducing patch sizes. To further study the reason for finding (3), we plot the performance of Varaint A (from ablation study A) under different patch sizes in Fig. 3a. We find that Variant A’s performance degrades when sequence length is more than $64$. This means that dynamic information selection based on edge-specific temporal patterns is essential for DyGMamba to optimally process long sequences.
@@
Figure 3: Impact of patch size on DyGFormer, DyGMamba and Varaint A, given a fixed number of sampled temporal neighbors $\rho$ on Enron. Patch size $p$ varies from $8$, $4$, $2$, $1$. Sequence length $\rho/p$ increases as patch size decreases. Performance is the transductive AP under random NSS.
@@
Figure 4: GPU memory consumption on Enron with increasing sampled neighbors. For clarity, memory consumption is shown only when neighbors are more than $512$.
@@
In historical interaction modeling, DyGMamba holds a computational complexity of $O(\rho/p)$ linear to $\rho/p$. While DyGFormer’s complexity is $O((\rho/p)^{2})$, which is quadratic to $\rho/p$. As a result, as the sequence length grows (either $\rho$ increases or $p$ decreases), DyGFormer is less scalable compared with DyGMamba. We present detailed computation and analysis of complexity in App. A.

### v2

@@ 3.2 Learning from Edge-Specific Temporal Patterns
To capture edge-specific temporal patterns, we use another time-level SSM block consisting of $l_{T}$ layers. We first find out $k$ temporally nearest historical interactions between $u$ and $v$ before $t$ and sort them in the chronological order, i.e., $\{(u,v,t_{0}),...,(u,v,t_{k-1})|t_{0}<...<t_{k-1}<t\}$. Then we construct a timestamp sequence $\{t_{0},t_{1},...,t_{k-1},t\}$ based on these interactions and the prediction timestamp $t$. We compute the time difference between each neighboring pair of them and further get a time difference sequence $\{t_{1}-t_{0},t_{2}-t_{1},...,t-t_{k-1}\}$, representing the change of time interval between two identical interactions. Each element in this sequence is input into the time encoding function stated above to get a edge-specific (specific to the edge $(u,v,t)$) time feature. The features are stacked into a feature matrix $\mathbf{H}_{u,v}^{t}\in\mathbb{R}^{k\times d_{T}}$ and mapped by an MLP $f_{\text{map}1}(\cdot):\mathbb{R}^{d_{T}}\to\mathbb{R}^{\gamma d}$ ($\gamma\in[0,1]$ is a hyperparameter), i.e., $\mathbf{H}_{u,v}^{t}:=f_{\text{map}1}(\mathbf{H}_{u,v}^{t})$. A time-level SSM layer takes $\mathbf{H}_{u,v}^{t}$ as input and computes
@@
After the time-level SSM block, we compute a compressed representation to represent the edge-specific temporal pattern by averaging over $k$ encoded time intervals: $\mathbf{h}_{u,v}^{t}=\text{MeanPooling}(\mathbf{H}_{u,v}^{t})$. As a result, we have $\mathbf{h}_{u,v}^{t}\in\mathbb{R}^{\gamma d}$ to represent the temporal pattern specific to the edge $(u,v,t)$. To leverage learned temporal pattern, we use it to dynamically select the information from the encoded temporal neighbors $\mathbf{H}_{\theta}^{t}$
$\hat{\mathbf{h}}_{u,v}^{t}=f_{\text{map}2}(\mathbf{h}_{u,v}^{t})\in\mathbb{R}^{4d};$ (7a)
$\hat{\mathbf{h}}_{\theta}^{t}={\mathbf{w}_{\text{agg}}}^{\top}\mathbf{H}_{\theta}^{t}\in\mathbb{R}^{4d},\ \text{where}\ \mathbf{w}_{\text{agg}}=f_{\text{map}3}(\mathbf{H}_{\theta}^{t})\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil};$ (7b)
$\alpha_{u}=f^{\prime}(\hat{\mathbf{h}}_{v}^{t})*\hat{\mathbf{h}}_{u,v}^{t}\in\mathbb{R}^{4d},\ \alpha_{v}=f^{\prime}(\hat{\mathbf{h}}_{u}^{t})*\hat{\mathbf{h}}_{u,v}^{t}\in\mathbb{R}^{4d};$ (7c)
$\mathbf{h}_{\theta}^{t}={\beta_{\theta}}^{\top}\mathbf{H}_{\theta}^{t},\ \text{where}\ \beta_{\theta}=\text{Softmax}(\mathbf{H}_{\theta}^{t}\alpha_{\theta})\in\mathbb{R}^{\lceil(|\textit{Nei}_{\theta}^{t}|+1)/p\rceil}.$ (7d)
$f_{\text{map}2}(\cdot):\mathbb{R}^{\gamma d}\to\mathbb{R}^{4d}$ and $f_{\text{map}3}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{1}$ are two mapping MLPs. $f^{\prime}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{4d}$ is another MLP introducing training parameters. Note that $\alpha_{u}$/$\alpha_{v}$ is computed by considering both the edge-specific temporal pattern and the opposite node $v$/$u$. In the node-level SSM block, we separately model the one-hop temporal neighbors of each node $\theta$, making it hard to connect $u$ and $v$. Computing $\alpha_{\theta}$ as Eq. 7c helps to strengthen the connection between both nodes and meanwhile incorporates the learned temporal pattern. $\beta_{\theta}$ is derived by transforming the queried results based on $\alpha_{\theta}$ into weights. It is then used to compute a weighted-sum of all temporal neighbors for representing $\theta$ at $t$, i.e., $\mathbf{h}_{\theta}^{t}$. Finally, we output the representations of $u$, $v$ and the edge-specific temporal pattern by employing two output MLPs $f_{\text{out}1}(\cdot):\mathbb{R}^{4d}\to\mathbb{R}^{d_{N}}$ and $f_{\text{out}2}(\cdot):\mathbb{R}^{\gamma d}\to\mathbb{R}^{d_{N}}$
@@ 3.3 Leveraging Learned Representations for Link Prediction
We leverage $\mathbf{h}_{\theta}^{t}$ and $\mathbf{h}_{u,v}^{t}$ for dynamic link prediction. We employ a prediction MLP, i.e., $f_{\text{LP}}(\cdot):\mathbb{R}^{3d_{N}}\to\mathbb{R}$, as the predictor. The probability of existing a link $(u,v,t)$ is computed as $y^{\prime}(u,v,t)=\text{Sigmoid}(f_{\text{LP}}(\mathbf{h}_{u}^{t}\|\mathbf{h}_{v}^{t}\|\mathbf{h}_{u,v}^{t}))$. For model parameter learning, we use the following loss function
$\mathcal{L}=-\frac{1}{2M}\sum_{2M}\left(y(u,v,t)\log(y^{\prime}(u,v,t))+(1-y(u,v,t))\log(1-y^{\prime}(u,v,t))\right).$ (9)
@@ 4 Experiments
In Sec. 4.2, we validate DyGMamba’s ability in CTDG representation learning by comparing it with baseline methods on dynamic link prediction [^2] To supplement, we also validate on the dynamic node classification task. Since current mainstream datasets of this task requires no long-term temporal reasoning, we put the discussion in App. F. We do further analysis to show the effectiveness of model components, in particular the time-level SSM block. In Sec. 4.3, we further show DyGMamba’s efficiency. We also show that it achieves much stronger scalability in modeling long-term temporal information compared with DyGFormer.
@@ 4.1 Experimental Setting
We consider seven real-world CTDG datasets collected by (Poursafaei et al., 2022), i.e., LastFM, Enron, MOOC, Reddit, Wikipedia, UCI and Social Evo.. Dataset statistics are presented in App. A.1. Among them, we take LastFM, Enron and MOOC as long-range temporal dependent datasets because according to Yu et al. (2023), much longer histories are needed for optimal representation learning on them. We compare DyGMamba with ten recent CTDG baseline models, i.e., JODIE (Kumar et al., 2019), DyRep (Trivedi et al., 2019), TGAT (Xu et al., 2020), TGN (Rossi et al., 2020), CAWN (Wang et al., 2021c), EdgeBank (Poursafaei et al., 2022), TCL (Wang et al., 2021a), GraphMixer (Cong et al., 2023), DyGFormer (Yu et al., 2023) and CTAN (Gravina et al., 2024). Among them, only DyGFormer and CTAN are designed for long-range temporal information propagation. Detailed descriptions of baseline methods are presented in App. B. We also implemented FreeDyG (Tian et al., 2024) by using its official code repository, however, on LastFM, we find that FreeDyG’s loss cannot converge and the reported results are not reproducible. So we do not report its performance in our paper.
Implementation Details and Evaluation Settings.
We use the implementations and the best hyperparameters provided by Yu et al. (2023) for all baseline models except CTAN. For CTAN, we use its official implementation, fixing the number of layers to 5. All models are trained with a batch size of 200 for fair efficiency analysis. For DyGMamba, we report the number of sampled one-hop temporal neighbors $\rho$ and the patch size $p$ here. On Wikipedia, Social Evo., and UCI, $\rho\ \&\ p=$ 32 & 1. On Reddit, $\rho\ \&\ p=$ 64 & 2. On MOOC, $\rho\ \&\ p$ = 128 & 4. On Enron, $\rho\ \&\ p=$256 & 8. On LastFM, $\rho\ \&\ p=$ 512 & 16. Note that to fairly compare DyGMamba’s efficiency with DyGFormer, we keep the sequence length $\rho/p$ input into the SSM as same as the length input into Transformer in Yu et al. (2023), i.e., $\rho/p=$ 32. All experiments are implemented with PyTorch (Paszke et al., 2019) on a server equipped with an AMD EPYC 7513 32-Core Processor and a single NVIDIA A40 with 45GB memory. We run each experiment for five times with five random seeds and report the mean results together with error bars. Further implementation details including complete hyperparamter configurations are presented in App. C. We employ two evaluation settings following previous works: the transductive and inductive settings. As suggested in (Poursafaei et al., 2022), we do link prediction evaluation using three negative sampling strategies (NSSs): random, historical and inductive. Historical NSS is only considered under the transductive setting. See App. D for detailed explanations. We employ two metrics, i.e., average precision (AP) and area under the receiver operating characteristic curve (AUC-ROC)
@@
4.2 Performance Analysis
Comparative Study on Benchmark Datasets.
We report the AP of transductive and inductive link prediction in Table 1 and 2 (AUC-ROC reported in Table 12 and 13 in App. E). We find that: (1) DyGMamba constantly ranks top $1$ under the random NSS, showing a superior performance; (2) Under the historical and inductive NSS, DyGMamba can achieve the best average rank compared with all baselines. More importantly, it shows more superiority on the datasets where encoding longer-term temporal dependencies is necessary, e.g., on LastFM, Enron and MOOC. (3) Among the models that can do long range propagation of information over time (i.e., DyGFormer, CTAN and DyGMamba), DyGMamba achieves the best average rank under any NSS setting in both transductive and inductive link prediction. On the long-range temporal dependent datasets, DyGMamba outperforms DyGFormer and CTAN in most cases; (4) CTAN achieves much better results in transductive than in inductive link prediction. This is because CTAN requires multi-hop temporal neighbors to learn node representations, which is difficult for unseen nodes. By contrast, DyGMamba and DyGFormer require only one-hop temporal neighbors, thus performing much better in inductive link prediction.
Table 3: Ablation studies under transductive setting. R/H/I means random/historical/inductive NSS. Metric is AP.
@@ 4.2 Performance Analysis
Table 4: Ablation studies under inductive setting. R/I means random/inductive NSS. Metric is AP.
@@
Ablation Study.
We conduct two ablation studies to study the effectiveness of model components. In study A, we make a model variant (Variant A) by removing the time-level SSM block and restrain our model from learning temporal patterns (information selection is substituted by mean pooling over the output of Eq. 5). In study B, we make a model variant (Variant B) by removing the Mamba SSM layers (Eq. 4) in the node-level SSM block. We find that: (1) From Table 3 and 4, we observe that under any NSS setting, Variant A is constantly beaten by the complete model, proving the effectiveness of learning edge-specific temporal patterns; (2) We also witness a performance drop from Table 3 and 4 if we switch DyGMamba to Variant B. This indicates the importance of encoding the one-hop temporal neighbors with SSM layers for capturing graph dynamics.
A Closer Look into Temporal Pattern Modeling.
We observe from ablation studies that time-level SSM block contributes to better model performance on real-world datasets. To better quantify the benefits brought by temporal pattern modeling, we construct three synthetic datasets, i.e., S1, S2 and S3, that follow different patterns and compare our model with DyGFormer, CTAN as well as Variant A on them. Each synthetic dataset contains 7 nodes, where the interactions of each pair of two nodes follow a certain pattern along time. And for each node, we generate interactions with all the other nodes. Assume we have a pair of node $u$ and $v$ and they have interactions at $\{t_{i}\}_{i=0}^{N}$, in S1, the time intervals between neighboring interactions $\{t_{1}-t_{0},...,t_{N}-t_{N-1}\}$ follow an increasing trend with a constant velocity of 0.05, i.e., $(t_{i+2}-t_{i+1})-(t_{i+1}-t_{i})=$ 0.05 . In S2, we set the time intervals to a decreasing trend with the same velocity, i.e., $(t_{i+1}-t_{i})-(t_{i+2}-t_{i+1})=$ 0.05. And in S3, we modify S1 by repeating several periods of increasing patterns taken from S1 to form a periodic dataset. In this way, we have three datasets demonstrating diverse temporal patterns: increasing/decreasing/periodic time intervals between neighboring interactions. Details of dataset construction and statistics are provided in App. A.2. From Table 5, we observe that DyGMamba greatly outperforms DyGFormer and CTAN. More importantly, Variant A shows similar performance to DyGFormer, meaning that our time-level SSM block is able to capture temporal patterns and modeling such patterns is important in CTDG reasoning. For more implementation details on synthetic datasets, please refer to App. C.2.
Table 5: Performance (Random NSS) on synthetic datasets.
Datasets DyGFormer CTAN Variant A DyGMamba
S1 55.19 $\pm$ 0.98 51.25 $\pm$ 2.11 53.72 $\pm$ 0.04 81.58 $\pm$ 1.31
S2 57.80 $\pm$ 4.61 51.17 $\pm$ 0.93 60.16 $\pm$ 2.20 85.36 $\pm$ 2.55
S3 79.20 $\pm$ 0.60 51.46 $\pm$ 0.19 76.36 $\pm$ 0.45 86.59 $\pm$ 0.09
(a) AP on synthetic datasets.
Datasets DyGFormer CTAN Variant A DyGMamba
S1 56.27 $\pm$ 0.54 51.25 $\pm$ 2.38 53.16 $\pm$ 0.39 86.61 $\pm$ 1.30
S2 59.06 $\pm$ 6.07 51.46 $\pm$ 0.91 62.50 $\pm$ 2.28 89.94 $\pm$ 2.70
S3 82.89 $\pm$ 1.34 52.12 $\pm$ 0.44 81.78 $\pm$ 0.40 91.72 $\pm$ 0.11
(b) AUC-ROC on synthetic datasets.
@@ 4.3 Efficiency Analysis
Figure 2: Efficiency comparison on four datasets among DyGMamba and five baselines in terms of number (#) of parameters, training time per epoch and GPU memory. The performance metric here is AP of transductive link prediction under random NSS. The greener, the better performance/efficiency. In contrast to other methods, DyGMamba consistently shows strong overall capability across different datasets. More explanations in Sec. 4.3.
@@
Fig. 2 compares DyGMamba with five baselines in terms of number of parameters (model size), per epoch training time and GPU memory consumption during training [^3] The baselines not included here are either extremely inefficient (e.g., CAWN) or inferior in performance (e.g., DyRep). Complete statistics of all baseline models presented in App. G. We find that: (1) DyGMamba uses very few parameters while maintaining the best performance, showing a strong parameter efficiency. Only CTAN constantly uses fewer parameters than DyGMamba, however, its performance is much worse; (2) DyGMamba is much more efficient than DyGFormer with the same length of input sequence ($\rho/p=$32); (3) Although DyGMamba generally consumes more GPU memory and per epoch training time compared with most baselines, the gap of consumption is not very large in most cases. To model more temporal neighbors for long-range temporal dependent datasets, DyGMamba naturally requires more computational resources, thus enlarging the consumption gap. DyGFormer shows the same trend as DyGMamba since it also captures long-term temporal dependencies by considering more temporal neighbors; (4) CTAN requires very few computational resources. However, on long-range temporal dependent datasets, it is beaten by DyGFormer and DyGMamba with a great margin. Besides, CTAN is also hard to converge. Although it requires very little per epoch training time, it requires much more epochs to reach the best performance, leading to a long total training time. See App. G.2 and G.3 for details.
Impact of Patch Size on Scalability and Performance.
@@
Patching treats $p$ temporal neighbors as one patch and thus decreases the sequence length by $p$ times. This is very helpful in cutting the consumption of GPU memory and training/evaluation time. However, patching introduces excessive parameters because it is done through $f_{N}$, $f_{E}$, $f_{T}$ and $f_{F}$ (Eq. 3) whose sizes increase as the patch size grows. Fig. 3b shows the numbers of parameters of DyGFormer and DyGmamba with different patch sizes on a long-term temporal dependent dataset Enron. We find that patching greatly affects model sizes. To further study how patching affects DyGMamba, we decrease the patch size gradually from 8 to 1 and track DyGMamba’s performance (Fig. 3a) as well as efficiency (Fig. 3b to 3d) on Enron. Meanwhile, we also keep track on DyGFormer under the same patch size for comparison. We have several findings: (1) Whatever the patch size is, DyGMamba always consumes fewer parameters, less GPU memory and per epoch training time, showing its high efficiency; (2) While both models require increasing computational budgets as the patch size decreases, the speed of increase is much lower for DyGMamba, demonstrating its strong scalability in modeling longer sequences; (3) Different trends in performance change are observed between two models. While DyGFormer performs worse, DyGMamba can benefit from a smaller patch size, indicating its strong ability to capture nuanced temporal details even if the sequence becomes much longer. Note that the models use fewer parameters under smaller patch size. This also shows that DyGMamba can achieve much stronger parameter efficiency by reducing patch sizes. To further study the reason for finding (3), we plot the performance of Variant A (from ablation study A) under different patch sizes in Fig. 3a. We find that Variant A’s performance degrades when sequence length is more than $64$. This means that dynamic information selection based on edge-specific temporal patterns is essential for DyGMamba to optimally process long sequences.
@@
Figure 3: Impact of patch size on DyGFormer, DyGMamba and Variant A, given a fixed number of sampled temporal neighbors $\rho$ on Enron. Patch size $p$ varies from 8, 4, 2, 1. Sequence length $\rho/p$ increases as patch size decreases. Performance is the transductive AP under random NSS.
@@
Figure 4: GPU memory consumption on Enron with increasing neighbors. For clarity, memory is shown only when neighbors are more than 512.
@@
In historical interaction modeling, DyGMamba holds a computational complexity linear to $\rho/p$, while DyGFormer’s complexity is quadratic to $\rho/p$. As a result, as the sequence length grows (either $\rho$ increases or $p$ decreases), DyGFormer is less scalable compared with DyGMamba.